Skip to main content
Thermo Fisher Scientific

Monochromator shift frozen

Issue

Mono shift X is not changing (V1x, V2x are stuck at –225V and 225V)

Environment

Themis Z

Resolution

These are possible resolutions for the issue:

  • Power restart of the FEG & HT tank could work, otherwise a patch is needed, contact GTS for "mono_dac_reinit.cmd"

Cause

These are possible causes for the issue:

  • After XFEG install and replacement of GMS, DPA, SDC, the high

Details

After replacing SDC, DPA, GMS and the XFEG for another issue, it was not possible to align the monoshift. It was diagnosed that the V1x and V2x were stuck at –225V and 225V.

The reason was not a mistake done during the install of the new XFEG and electronic boards. It was simply a software bug which sets the SDC2 high-res DACs from the Mono to a max value in X which cannot get recovered.

What's happening, is the high-resolution DAC on SDC board goes into 'reset state' and cannot recover. If the issue is with the SDC high resolution DAC that is getting into a reset mode, power restart of the FEG & HT tank could also work to recover it. Alternatively, a patch from GTS is needed "mono_dac_reinit.cmd".

Patch:
Attached software patch.

In case mono freezing issue occurs again: In case mono freezing issue occurs again:

  1. Open TEM HAL
  2. Load CPBm
  3. Adjust voltage of mainplate1 and mainplate2 manually. These control the mono shift x. See if the beam reacts.
  4. Close TEM server.
  5. Run mono_dac_reinit.cmd
  6. Restart TEM server.

Before patch:
clipboard_efd1b97a4a04b27e8259ae5f95af40764.png

After patch:
clipboard_efc0ee72aae620dc7d1ad1763f500d176.png

 

  • Was this article helpful?